OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half.

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x- x1) when given a point (x1, y1) and m = slope so plug in what you know the point (6, −3) and has a slope of 1/2 so plug in x1 = 6 , y1 = -3 , m = 1/2

OpenStudy (anonymous):

Thank you

jigglypuff314 (jigglypuff314):

your welcome :)